# Fair Foundation Models for Medical Image Analysis: Challenges and Perspectives

### v1

@@ 4.3 Deployment
Documentation. Comprehensive documentation constitutes a fundamental requirement for the ethical implementation of FMs 11, 127. The documentation must encompass detailed specifications of training data sources, testing benchmark methodologies, and explicit deployment guidelines that minimize risk and bias 128. Moreover, the documentation should present information in a stratified manner, ensuring accessibility across varying levels of technical expertise while maintaining rigorous detail. This multi-level documentation approach enables stakeholders to comprehend the model’s inherent risks, biases, and operational constraints. Such transparency serves as a crucial mechanism for fostering trust and promoting responsible model deployment in clinical settings.
License. Open FMs represent a strategic approach to addressing global bias mitigation, enhancing transparency, and facilitating equitable power distribution 129, 130. By providing unrestricted access to data, code, and model weights, this approach enables resource-constrained institutions to engage in model development and adaptation, thereby fostering innovation through reduced computational barriers. As established in Section 4.1.2, while closed models present substantial impediments to achieving fairness and impartial outcomes, open-weight architectures facilitate superior downstream task adaptation. Nevertheless, ensuring adherence to intended model applications remains crucial for preventing bias propagation and maintaining equity. The implementation of a Responsible AI License framework provides a structured mechanism for guiding ethical and equitable model utilization.
Monitoring. In medical settings, the prevalence of the disease and the distribution of patients accessing a specific hospital can change over time. While assessing fairness and bias during pretraining and fine-tuning is critical for FMs, continuous monitoring of deployed models in real-world scenarios is equally important. Despite its importance, such monitoring is not widely practiced; in the United States, only 44% of institutions reported conducting local evaluations for bias 131.
@@ 5 Policymakers
This section examines two critical aspects of FMs development: governance frameworks for ensuring reliability and the strategic allocation of essential resources.
@@ 5.1 Governance
Reliability. The implementation of FMs in healthcare faces critical reliability challenges due to insufficient standardization in bias detection and mitigation practices 19, 20, 21, 25, 114. Despite a growing emphasis on AI ethics, healthcare institutions lack both established evaluation frameworks and qualified experts needed to assess these systems effectively. This deficiency is particularly concerning because bias continues to emerge in deployment scenarios, while current mitigation methods demonstrate limited reliability 25, 114. The challenge of bias mitigation becomes particularly critical in regions with limited data representation, notably in Global South nations (Figure 2), where healthcare systems encounter substantial obstacles in implementing and adapting these models for their populations. This disparity underscores a critical gap in current FMs deployment strategies, as the absence of representative data further compounds existing healthcare inequities.
FMs as institutions. Research demonstrates that algorithmic systems function as institutional frameworks that organize complex machine-human interactions in decision-making processes 132. FMs represent a significant evolution in this infrastructure, enabling complex decision-making capabilities across medical domains, offering pre-trained architectures adaptable to diverse medical tasks. However, these models systematically perpetuate societal power imbalances and biases through their operational frameworks 133. Conceptualizing FMs as institutions, rather than purely technical implementations, facilitates the establishment of comprehensive evaluation protocols and governance frameworks that enhance model reliability through systematic oversight, standardized assessment criteria, and robust accountability measures.
Regulation. Legislative frameworks are crucial in enhancing AI system reliability in healthcare. The EU AI Act exemplifies a pioneering approach by classifying medical AI systems as "high-risk" and establishing provisions for General Purpose AI Models (GPAI), which encompass FMs. These models present distinct challenges due to their scale, adaptability, and cross-domain impact 134. While such oversight aims to enhance safety and accountability, implementation barriers emerge 135. Studies reveal that companies struggle with extensive technical documentation requirements and associated costs 136. These obstacles become more pronounced in developing nations, where compliance expenses may hinder local innovation. Although regulations could ensure unbiased systems and improve data practices, stringent requirements risk impeding AI advancement in emerging economies. This creates tension between maintaining safety standards and fostering equitable technological progress globally.
@@ 5.2 Resource Allocation
Workforce. Fair FMs demand interdisciplinary collaboration among technical, legal, and social experts to ensure comprehensive bias mitigation 13, 137. These multidisciplinary teams integrate diverse societal values and ethical perspectives into model development and evaluation protocols. Essential to this process is the active participation of marginalized communities, whose insights prove crucial for identifying systemic biases in FMs. The global distribution of FM expertise, however, remains heavily concentrated in the United States and China 138, 139, creating significant barriers to diverse team formation. This geographic concentration is exacerbated by linguistic constraints: the predominance of English in educational materials and technical documentation impedes knowledge dissemination and FM expertise development in non-English-speaking regions.
Data source. AI research disparities manifest critically in data infrastructure and computational resources across global regions. Figure 2 illustrates the significant underrepresentation of Global South populations in major medical imaging datasets. Quantitative analyses reveal systematic exclusion across modalities, with Global South regions constituting less than 0.7% of text-domain datasets 140. This limitation extends to linguistic diversity: among medical imaging repositories, only PadChest provides non-English (Spanish) annotations (Table 1). Developing equitable FMs requires datasets that reflect global population diversity across both geographic and demographic dimensions. Achieving representative data collection necessitates integrated technical and governance frameworks that actively incorporate marginalized populations into healthcare data systems 32, 33, 34, 35.
Computing. FMs development requires extensive computational infrastructure, limiting access to organizations with advanced resources 133. Analysis of the November 2024 Top500 supercomputer rankings demonstrates pronounced regional disparities: North America holds 55.6% of capacity, Europe 27.4%, and Asia 15.9%, while South America and Oceania represent just 0.6% and 0.5% respectively. The combined scarcity of computational power and large-scale datasets creates a fundamental barrier in the Global South, where only Brazil, Australia, and Argentina maintain Top500-listed facilities. Africa’s absence from these rankings underscores how infrastructure limitations constrain both FMs development and deployment.
@@ 6 Conclusion
FMs represent a transformative advancement in medical imaging analysis, yet their implementation presents both opportunities and challenges for achieving equitable healthcare delivery. Our comprehensive review demonstrates that effective bias mitigation in FMs requires systematic interventions throughout the development pipeline, from data curation to deployment protocols. While technical innovations in training methodologies show promise for enhancing fairness without relying on sensitive attributes, the substantial computational and data demands of these models risk exacerbating global inequalities. The emergence of regulatory frameworks such as the EU AI Act reflects growing recognition of FMs societal impact and the need for governance structures that ensure responsible development. However, significant disparities persist, particularly in Global South nations where limited access to essential resources including specialized workforce, datasets, and computational infrastructure hinders both development and implementation of fair FMs. Moving forward, addressing these challenges requires coordinated action between technologists, healthcare providers, and policymakers to develop accessible solutions and appropriate frameworks for low-resource countries and institutions. As FMs continue to evolve, their successful implementation in healthcare will depend on our ability to balance technical innovation with ethical principles, ultimately working toward reducing rather than amplifying existing healthcare disparities.
@@ Acknowledgements
We thank Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP), grants 21/14725-3 and 23/12493-3, Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPQ), Swiss National Science Foundation (SNSF) under Grant No. 200021E_214653, Santos Dumont supercomputer at the LNCC.

### v2

@@ 4.3 Deployment
Documentation. Comprehensive documentation constitutes a fundamental requirement for the ethical implementation of FMs [94, 18]. The documentation must encompass detailed specifications of training data sources, testing benchmark methodologies, and explicit deployment guidelines that minimize risk and bias [109]. Moreover, the documentation should present information in a stratified manner, ensuring accessibility across varying levels of technical expertise while maintaining rigorous detail. This multi-level documentation approach enables stakeholders to comprehend the model’s inherent risks, biases, and operational constraints. Such transparency serves as a crucial mechanism for fostering trust and promoting responsible model deployment in clinical settings.
License. Open FMs represent a strategic approach to addressing global bias mitigation, enhancing transparency, and facilitating equitable power distribution [17, 103]. By providing unrestricted access to data, code, and model weights, this approach enables resource-constrained institutions to engage in model development and adaptation, thereby fostering innovation through reduced computational barriers. As established in Section 4.1.2, while closed models present substantial impediments to achieving fairness and impartial outcomes, open-weight architectures facilitate superior downstream task adaptation. Nevertheless, ensuring adherence to intended model applications remains crucial for preventing bias propagation and maintaining equity. The implementation of a Responsible AI License framework provides a structured mechanism for guiding ethical and equitable model utilization.
Monitoring. In medical settings, the prevalence of the disease and the distribution of patients accessing a specific hospital can change over time. While assessing fairness and bias during pretraining and fine-tuning is critical for FMs, continuous monitoring of deployed models in real-world scenarios is equally important. Despite its importance, such monitoring is not widely practiced; in the United States, only 44% of institutions reported conducting local evaluations for bias [114].
@@ 5 Policymakers
This section examines two critical aspects of FMs development: governance frameworks for ensuring reliability and the strategic allocation of essential resources. Figure 3 summarizes the key points discussed in the text, providing a comprehensive overview of foundation model development along with the cost and importance of each step.
Figure 3: Overview recommendation: The figure summarizes recommendations for achieving fair foundation models, indicating for each stage its medical infrastructure, importance for fairness (essential), high cost, and computation. The recommendations focus on where governments and industry should prioritize technical interventions to advance fairness. Medical infrastructure is critical for both downstream applications and data creation, but their cost profiles differ: data creation depends primarily on expensive imaging equipment and large-scale data storage, whereas downstream tasks require substantial clinical expertise for data annotation. With respect to fairness mitigation, the most critical and challenging stage is pre-training, which demands extensive computational infrastructure and incurs high costs. The figure also highlights evaluation and data curation as comparatively low-cost yet important strategies for improving fairness. Generative models offer a promising avenue to balance data creation and alleviate infrastructure constraints by better representing the underlying population. At the same time, real-data collection remains more critical for fairness, because equitable access to medical infrastructure across all population groups is essential, whereas generative models can only complement these efforts by improving population representation in the training data. Finally, agglomerative models and world models are identified as future directions with potential to enhance fairness, although they are tightly coupled to pre-training and therefore are not quantitatively assessed in the figure.
@@ 5.1 Governance
Reliability. The implementation of FMs in healthcare faces critical reliability challenges due to insufficient standardization in bias and hallucinations detection and mitigation practices [77, 53, 88, 71, 176]. Despite a growing emphasis on AI ethics, healthcare institutions lack both established evaluation frameworks and qualified experts needed to assess these systems effectively. This deficiency is particularly concerning because bias continues to emerge in deployment scenarios, while current mitigation methods demonstrate limited reliability [71, 176]. The challenge of bias mitigation becomes particularly critical in regions with limited data representation, notably in Global South nations (Figure 2), where healthcare systems encounter substantial obstacles in implementing and adapting these models for their populations. This disparity underscores a critical gap in current FMs deployment strategies, as the absence of representative data further compounds existing healthcare inequities.
FMs as institutions. Research demonstrates that algorithmic systems function as institutional frameworks that organize complex machine-human interactions in decision-making processes [4]. FMs represent a significant evolution in this infrastructure, enabling complex decision-making capabilities across medical domains, offering pre-trained architectures adaptable to diverse medical tasks. However, these models systematically perpetuate societal power imbalances and biases through their operational frameworks [35]. Conceptualizing FMs as institutions, rather than purely technical implementations, facilitates the establishment of comprehensive evaluation protocols and governance frameworks that enhance model reliability through systematic oversight, standardized assessment criteria, and robust accountability measures.
Regulation. Legislative frameworks are crucial in enhancing AI system reliability in healthcare. The EU AI Act exemplifies a pioneering approach by classifying medical AI systems as "high-risk" and establishing provisions for General Purpose AI Models (GPAI), which encompass FMs. These models present distinct challenges due to their scale, adaptability, and cross-domain impact [108]. The EU AI Act mandates requirements to detect, prevent, and mitigate bias in applications that could result in discrimination. These regulatory requirements transform FMs development pipelines by necessitating the adoption of fairness-enhancing techniques discussed in this work. Although these changes impose extensive technical documentation requirements and associated costs [1, 24], the adoption of such pipelines enhances the reliability of companie’s and government products and enables the scalable deployment of AI from local settings to national and global levels.
@@ 5.2 Resource Allocation
Workforce. Fair FMs demand interdisciplinary collaboration among technical, legal, and social experts to ensure comprehensive bias mitigation [85, 14]. These multidisciplinary teams integrate diverse societal values and ethical perspectives into model development and evaluation protocols. Essential to this process is the active participation of marginalized communities, whose insights prove crucial for identifying systemic biases in FMs. The global distribution of FM expertise, however, remains heavily concentrated in the United States and China [5, 178], creating significant barriers to diverse team formation. Another challenge arises from data sources and computing resources; many countries face difficulties in providing high-performance computing (HPC) access to institutions. A valuable contribution in this regard is the establishment of training centers that offer education, data, and computing resources.
Data source. AI research disparities manifest critically in data infrastructure and computational resources across global regions. Figure 2 illustrates the significant underrepresentation of Global South populations in major medical imaging datasets. Quantitative analyses reveal systematic exclusion across modalities, with Global South regions constituting less than 0.7% of text-domain datasets [95]. This limitation extends to linguistic diversity: among medical imaging repositories, only PadChest provides non-English (Spanish) annotations (Table 2). Developing equitable FMs requires datasets that reflect global population diversity across both geographic and demographic dimensions. Achieving representative data collection necessitates integrated technical and governance frameworks that actively incorporate marginalized populations into healthcare data systems [154, 28, 11, 152].
Computing. FMs development requires extensive computational infrastructure, limiting access to organizations with advanced resources [35]. Analysis of the November 2024 Top500 supercomputer rankings demonstrates pronounced regional disparities: North America holds 55.6% of capacity, Europe 27.4%, and Asia 15.9%, while South America and Oceania represent just 0.6% and 0.5% respectively. The combined scarcity of computational power and large-scale datasets creates a fundamental barrier in the Global South, where only Brazil, Australia, and Argentina maintain Top500-listed facilities. Africa’s absence from these rankings underscores how infrastructure limitations constrain both FMs development and deployment.
@@ 6 Conclusion
FMs represent a transformative advancement in medical imaging analysis, yet their implementation presents both opportunities and challenges for achieving equitable healthcare delivery. Our comprehensive review demonstrates that effective bias mitigation in FMs requires systematic interventions throughout the development pipeline, from data curation to deployment protocols. While technical innovations in training methodologies show promise for enhancing fairness without relying on protected attributes, the substantial computational and data demands of these models risk exacerbating global inequalities. The emergence of regulatory frameworks such as the EU AI Act reflects growing recognition of FMs societal impact and the need for governance structures that ensure responsible development. However, significant disparities persist, particularly in Global South nations where limited access to essential resources including specialized workforce, datasets, and computational infrastructure hinders both development and implementation of fair FMs. Moving forward, addressing these challenges requires coordinated action between technologists, healthcare providers, and policymakers to develop accessible solutions and appropriate frameworks for low-resource countries and institutions. As FMs continue to evolve, their successful implementation in healthcare will depend on our ability to balance technical innovation with ethical principles, ultimately working toward reducing rather than amplifying existing healthcare disparities.
@@ Acknowledgements
We thank Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP), grants 21/14725-3 and 23/12493-3, Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPQ), Swiss National Science Foundation (SNSF) under Grant No. 200021E_214653, Santos Dumont supercomputer at the LNCC. We would like thank Lucas Tosta for the designs.